### v1

# Choi-Defined Resource Theories

Abstract
The resource theories of separable entanglement, non-positive partial transpose entanglement, magic, and imaginarity share an interesting property: An operation is free if and only if its renormalized Choi matrix is a free state. In this letter, we refer to resource theories exhibiting this property as Choi-defined resource theories. We demonstrate how and under what conditions one can construct a Choi-defined resource theory, and we prove that when such a construction is possible, the free operations are all and only the completely resource non-generating operations.
§ I.  Introduction
The recent booming interest in quantum computing, communication, and information originates from the understanding that quantumness allows us to outperform the classical world, which means that quantum objects are resources. The new technological developments, which are now becoming commercially available, demand a rigorous and thorough understanding of quantum resources, in order to quantify them precisely, and optimize their use. This is accomplished with the framework of resource theories [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. Resource theories are based on the simple observation that agents in a laboratory can only perform a subset of all possible deterministic quantum operations due to the various constraints on the agents and the systems involved. For instance, when two agents, Alice and Bob, are physically separated, they may not have a reliable means to exchange quantum systems, limiting them to local operations and classical communication (LOCC) [5, 6, 7]. Such allowed operations are called free. While free states are typically defined as those states that can be prepared with free operations, there are situations where it is beneficial to construct free operations starting from a set of free states. As an example, Alice and Bob may quickly realize that they can create all and only separable states and wonder what free operations are compatible with their set of free states. There are different ways to construct free operations starting from free states [1]. However, the minimal condition is that one cannot generate a resource by applying a free operation on a free state in a complete sense, i.e., even if applied only to part of a bipartite state. Operations defined with such a construction are known as completely resource non-generating (CRNG) operations [1]. CRNG operations are of particular interest because they are the largest set of free operations compatible with a given set of free states, and they can be used to derive lower or upper bounds in any resource theory with the same set of free states. For example, separable and positive partial transpose operations have been used to overcome the challenging characterization of LOCC operations [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. These extensions of the resource theory of entanglement are known as the resource theories of separable entanglement (SEP) and non positive partial transpose entanglement (NPT), respectively. In other settings, such as in the resource theory of magic [28, 29, 30, 31, 32] and imaginarity [33, 34, 35] the CRNG operations are even more relevant because they coincide with the free operations [31, 32, 35]. Remarkably, one notices that in each of these instances of resource theories a channel is considered free if and only if its renormalized Choi matrix is a free state [37, 38, 46, 32, 35]. We call resource theories that exhibit this property Choi-defined resource theories. This is a highly valuable property: Every problem associated with free quantum channels can be converted into a problem involving free states, which, in turn, can be conveniently tackled with tools developed in linear algebra, matrix analysis, and convex optimization [37, 38, 32, 35]. Given the significance of Choi-defined resource theories and their computational convenience one wonders when it is possible to construct such resource theories from a set of free states. In this letter, we answer this question by providing necessary and sufficient conditions on the set of free states. Moreover, we show that when such a construction is possible, the resulting set of free operations coincides with the set of CRNG operations associated with those free states. These results generalize what has been already observed in the resource theories of magic, imaginarity, NPT, and SEP [37, 38, 46, 31, 32, 35].
§ II.  Preliminaries
A quantum resource theory provides a partition of all quantum channels into free and resourceful ones, with the following properties: [1, 2]
  1. 1.
    the identity channel is free,
  2. 2.
    the swap channel is free,
  3. 3.
    discarding a system is free,
  4. 4.
    sequential composition of free channels is free,
  5. 5.
    parallel composition of free channels is free
In other words, an agent is allowed to do nothing on a system, to change the order in which they describe systems, and to discard a system. In addition, if they perform free operations in sequence (Figure 1), or in parallel, the resulting operation is still free. As mentioned in the introduction, we say that a state is free if it can be prepared with free operations. Moreover, if one applies a free channel to a free state one obtains a free state, meaning that it is impossible to obtain a resource from free objects. This is sometimes called the golden rule of resource theories [1]. As a consequence, every set of free states satisfies the following properties: It is closed under tensor product, partial tracing, and system swapping.
Figure 1: Sequential composition of operations. A source emits a quantum system, e.g., photon (colored in red). A first operation $\mathcal{M}$ is performed on this quantum system and produces a new quantum system as output. A second operation $\mathcal{N}$ is performed on the new quantum system. The overall operation done on the first (red) quantum system is the sequential composition of $\mathcal{M}$ and $\mathcal{N}$, denoted with $\mathcal{N}\circ\mathcal{M}$. When dealing with Choi matrices, the sequential composition is translated into the link product.
Remark 1.
One of the constraints of the resource theory of bipartite entanglement is that the two physically separated agents cannot exchange quantum systems. This may appear to conflict with the condition that the swap channel is a free operation. However, in this scenario, the systems in question are bipartite systems [2], i.e., $A=A_{1}A_{2}$ and $B=B_{1}B_{2}$. Here the physical separation is between subscript $1$ and $2$. The swap channel simply swaps the order of the two bipartite systems, for example, $A_{1}A_{2}B_{1}B_{2}\leftrightarrow B_{1}B_{2}A_{1}A_{2}$. If we organize the systems according to the physical separation, the action of the swap channel is $A_{1}B_{1}A_{2}B_{2}\to B_{1}A_{1}B_{2}A_{2}$. Therefore, the two agents are locally exchanging the order of their own quantum systems and there is no exchange of quantum systems between them.
Free operations are mathematically described by quantum channels. A quantum channel $\mathcal{M}_{A\to B}$ can be mapped to a bipartite matrix $M_{B\mathord{:}A^{\prime}}$ by using the Choi isomorphism. The matrix $M_{B\mathord{:}A^{\prime}}$ is known as the Choi-matrix of $\mathcal{M}_{A\to B}$ and it is defined as
$M_{B\mathord{:}A^{\prime}}=\mathcal{M}_{A\to B}\otimes\mathcal{I}_{A^{\prime}}(\Phi_{AA^{\prime}}),$ (1)
[…]
$\mathcal{M}_{A\to B}(\rho_{A})$ $=\operatorname{Tr}_{A^{\prime}A}[(M_{B\mathord{:}A^{\prime}}\otimes\rho_{A})(\operatorname{\mathds{1}}_{B}\otimes\Phi_{A^{\prime}A})]$ (2)
$=\operatorname{Tr}_{A}[M_{B\mathord{:}A}(\operatorname{\mathds{1}}_{B}\otimes\rho_{A}^{T})].$
Note that here one needs to know which system is the input of the original channel and which is the output. Clearly, if such information is missing, a matrix $M_{BA}$ could be associated with different linear maps, see the Supplemental Material (SM) for examples. Observe that, if we divide the Choi matrix $M_{B\mathord{:}A^{\prime}}$ of the channel $\mathcal{M}_{A\to B}$ by the dimension $d_{A}$ of $A$, we obtain a quantum state $\mu_{B\mathord{:}A^{\prime}}$ such that
$\operatorname{Tr}_{B}\mu_{B\mathord{:}A^{\prime}}=\frac{1}{d_{A}}\operatorname{\mathds{1}}_{A^{\prime}}.$ (3)
As a consequence, a bipartite state is the renormalized Choi matrix of a quantum channel if it satisfies the condition in Eq. (3) .
§ III.  Choi-defined resource theories
As noted above, the resource theory of magic, imaginarity, SEP, and NPT share an interesting property: A channel is free if and only if its renormalized Choi matrix is a free state. Now we generalize such resource theories by introducing the construction of Choi-defined operations, and Choi-defined resource theories.
Definition 1.
The Choi-defined (CD) operations associated with a set of free states are all and only the quantum channels such that their renormalized Choi matrix is a free state.
With this definition at hand, we are ready to define the Choi-defined resource theories.
Definition 2.
A quantum resource theory is a Choi-defined resource theory (CDRT) if its free operations coincide with the Choi-defined operations associated with its set of free states.
Clearly, not all quantum resource theories are CDRTs. For example, in the resource theory of athermality [16, 17] there is only one free state per system and several free quantum channels, therefore there is no one-to-one correspondence between free states and free channels. A more interesting example is the resource theory of entanglement. In this case, free states are separable states, but the construction of CD operations produces all separable operations, which are a larger set than LOCC [49, 50]. As a result, the resource theory of LOCC entanglement is not a CDRT.
§ IV.  Construction of Choi-defined resource theories
It is a natural question to characterize when, given a set of free states, the construction of a CDRT is allowed. In every well-defined resource theory, the identity is a free operations. As a consequence, its renormalized Choi matrix $\frac{1}{d_{A}}\Phi_{AA^{\prime}}$ is free in every CDRT. Requesting that resource theories are closed under the sequential composition of free channels (Figure 1), we find a less trivial condition for the set free states. In Ref. [47, 48], the authors introduced an operation between bipartite matrices, called link product, defined as
$N_{C\mathord{:}B^{\prime}}*M_{B\mathord{:}A^{\prime}}$ (4)
$=\operatorname{Tr}_{BB^{\prime}}[(N_{C\mathord{:}B^{\prime}}\otimes M_{B\mathord{:}A^{\prime}})(\operatorname{\mathds{1}}_{C}\otimes\Phi_{B^{\prime}B}\otimes\operatorname{\mathds{1}}_{A^{\prime}})]$
$=\operatorname{Tr}_{B}[(N_{C\mathord{:}B}\otimes\operatorname{\mathds{1}}_{A^{\prime}})(\operatorname{\mathds{1}}_{C}\otimes M_{B\mathord{:}A}^{T_{B}})],$
where $\cdot^{T_{B}}$ is the partial transpose on system $B$. This operation translates sequential composition of channels into an operation between Choi matrices. Thanks to the link product, we can restate the golden rule of resource theories as a condition on the set of free states. That is, the condition that $\mathcal{M}_{A\to B}(\rho_{A})$ be free whenever $\mathcal{M}_{A\to B}$ and $\rho_{A}$ are free becomes $d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}$ is a free state if $\rho_{A}$ and $\mu_{B\mathord{:}A^{\prime}}$ are free states and $\mu_{B\mathord{:}A^{\prime}}$ is the renormalized Choi matrix of a quantum channel. We now state our first main result.
Theorem 3.
It is possible to construct a CDRT associated with a set of free states if and only if for all systems $A$, and $B$
  1. 1.
    the state $\frac{1}{d_{A}}\Phi_{AA^{\prime}}$ is free,
  2. 2.
    if $\rho_{A}$ and $\mu_{B\mathord{:}A^{\prime}}$ are free states, and $\mu_{B\mathord{:}A^{\prime}}$ is the renormalized Choi matrix of a quantum channel, then $d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}$ is a free state.
We provide a detailed proof of this theorem in the SM. From the same proof we obtain also the following corollary.
Corollary 4.
In every CDRT the maximally mixed state is free.
An interesting common property of the examples of CDRTs presented in the introduction is that their free operations coincide with the CRNG operations [37, 38, 46, 31, 32, 35]. As we show in the next theorem, this is a general feature of every CDRT.
Theorem 5.
In every CDRT, free operations coincide with CRNG operations.
Notably, this theorem provides an easy construction for CRNG operations when the set of free states satisfies the conditions of Theorem 3. For example, the set of separable states satisfies the conditions of Theorem 3, as we show later. Therefore, Theorems 3 and 5 imply that the CRNG operations are those that have a separable Choi matrix, i.e., the separable operations. Note that in general CRNG operations are not CD operations. Indeed, the CRNG operations in the resource theory of athermality are the Gibbs-preserving operations (see, e.g., Ref. [51]), but their renormalized Choi matrices are not free states.
§ V.  Example: resource theory of magic
Now we give a closer look at the resource theory of magic [28, 29, 30, 31, 32]. It originates from the observation that it is possible to achieve universal quantum computation by using only Clifford unitaries and some special states called magic states [52]. Being grounded in quantum computation, systems are usually systems of $n$ qubits [30]. In this scenario, Clifford unitaries, state preparation and measurements in the computational basis are easy to perform, and therefore free. In particular, Clifford unitaries are those transformations that can be implemented with circuits containing only Hadamard, CNOT, and phase-shift gates [53]. The free operations are called stabilizer operations, and analogously free states are called stabilizer states. The resources are magic states and operations.
We prove now that the set of magic states satisfies the conditions of Theorem 3. For the first condition, we denote the number of qubits in $A$ with $n$, and the elements of the computational basis with |$\Ket{x_{1}\dots x_{n}}$, where $x_{j}=0,1$. The normalized Choi state |$\Ket{\phi}=\frac{1}{2^{n/2}}\sum_{x_{1}\dots x_{n}}\Ket{x_{1}\dots x_{n}}\Ket{x_{1^{\prime}}\dots x_{n^{\prime}}}$ is obtained by preparing the $n$ entangled states $\frac{1}{\sqrt{2}}\sum_{x_{j}}\Ket{x_{j}x_{j^{\prime}}}$ and by swapping the order of the systems. The $j$-th 2-qubit entangled state can be obtained with an Hadamard and a CNOT gate acting on |$\Ket{0_{j}0_{j^{\prime}}}$, and the swap channels can be decomposed into three CNOT gates (see, e.g., Ref. [54]). All these operations are stabilizer operations, therefore the renormalized Choi state $\frac{1}{d_{A}}\Phi_{AA^{\prime}}$ is a stabilizer state. This proves that condition 1 of Theorem 3 is satisfied. We now consider the closure under link product of the set of stabilizer states. Let $\mu_{BA^{\prime}}$ and $\rho_{A}$ be free states, where $A$ is a $n$-qubit system and $B$ is a $m$-qubit system. This implies that there exists a protocol to create $\mu_{BA^{\prime}}\otimes{\rho_{A}}$ for free. By Eq. (3), the state $d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}$ is obtained by a protocol with postselection:
  1. 1.
    Prepare $\mu_{BA^{\prime}}\otimes\rho_{A}$.
  2. 2.
    Perform a measurement in the Bell basis for each of the pair of qubits in $A^{\prime}A$
  3. 3.
    Postselect on the measurement outcome associated with $\frac{1}{d_{A}}\Phi_{A^{\prime}A}$. After the postselection, the system $B$ is in the state $d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}$, as requested.
Each of the steps of this protocol is a free operation. Indeed, the measurement in the Bell basis can be converted into a measurement in the computational basis by adding Hadamard and CNOT gates and these are stabilizer operations. Therefore the state $d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}$ is free, as requested by Theorem 3. The set of stabilizer states satisfy all the conditions of Theorem 3, and the CDRT constructed form it is the resource theory of magic [32]. Note that, by Theorem 5, in the resource theory of magic the free operations coincide with the CRNG operations. Note that Corollary 4 allows us to predict that the maximally mixed $n$-qubit state $\frac{1}{2^{n}}\operatorname{\mathds{1}}_{2}\otimes\dots\otimes\operatorname{\mathds{1}}_{2}$ is free. We can see this directly from the definition of stabilizer states. Indeed, each of the qubits is prepared as a probabilistic mixture of elements of the computational basis, and any state preparation in the computational basis is a free operation.
§ VI.  Example: resource theory of imaginarity
We now turn our attention to the resource theory of imaginarity [33, 34, 35]. The aim of this resource theory is to evaluate the role of imaginarity in quantum mechanics. A state is free if its density matrix, expressed with respect to a fixed basis, is real. States with density matrices containing non-real numbers are resources. The Choi state $\frac{1}{d_{A}}\Phi_{AA^{\prime}}$ relative to the fixed basis, has only real coefficients, and therefore is free. This proves condition 1 of Theorem 3. Condition 2 follows trivially because it can be expressed using only multiplications and traces of real matrices, from which it is possible to obtain only real matrices. This proves that it is possible to construct a CDRT from the set of real matrices, and it coincides with the resource theory of imaginarity, where the free operations are all the CRNG operations [35]. Note that, the maximally mixed state is real too, and therefore free, as predicted by Corollary 4.
§ VII.  Example: resource theory of separable entanglement
In the theory of entanglement [5, 6, 7, 8, 9, 10] systems are bipartite systems, and free states are separable states. In other words, each of the systems $A$ and $B$ mentioned in Theorem 3 is bipartite, i.e., $A=A_{1}A_{2}$, $B=B_{1}B_{2}$, and the physical separation is between systems with ‘$1$’ as subscript and systems with ‘$2$’ as subscript. To prove the first condition of Theorem 3, we fix an ortonormal basis for each system, e.g., {|$\Set{\Ket{a_{1}}}$ is a basis for $A_{1}$, and we write the Choi state as
$\Phi_{AA^{\prime}}$ $=\sum_{a_{1},a_{2}}|a_{1}a_{2}\rangle\!\langle a^{\prime}_{1}a^{\prime}_{2}|$ (5)
$=\sum_{a_{1}}|a_{1}\rangle\!\langle a^{\prime}_{1}|\otimes\sum_{a_{2}}|a_{2}\rangle\!\langle a^{\prime}_{2}|.$
Therefore $\frac{1}{d_{A}}\Phi_{AA^{\prime}}$ is separable, free, and condition 1 of Theorem 3 is satisfied. We now prove that the set of separable state is closed under link product. Let $\mu_{BA^{\prime}}$ be a separable state, i.e., there exists density matrices $\mu^{j}_{B_{1}A^{\prime}_{1}}$ and $\tilde{\mu}^{j}_{B_{2}A^{\prime}_{2}}$ such that $\mu_{B_{1}A^{\prime}_{1}B_{2}A^{\prime}_{2}}=\sum_{j}p_{j}\mu^{j}_{B_{1}A^{\prime}_{1}}\otimes\tilde{\mu}^{j}_{B_{2}A^{\prime}_{2}}$, where $p_{j}\geq 0$ and $\sum_{j}p_{j}=1$. Let $\rho_{A}=\sum_{k}q_{k}\rho^{k}_{A_{1}}\otimes\tilde{\rho}^{k}_{A_{2}}$, where $q_{k}\geq 0$ and $\sum_{k}q_{k}=1$, be another separable state. Then
$\mu_{B\mathord{:}A^{\prime}}*\rho_{A}$ $=\operatorname{Tr}_{A}[\mu_{BA}(\operatorname{\mathds{1}}_{B}\otimes\rho_{A}^{T})]$ (6)
$=\sum_{j,k}p_{j}q_{k}\sigma^{j,k}_{B_{1}}\otimes\tilde{\sigma}^{j,k}_{B_{2}}.$
where $\sigma^{j,k}_{B_{1}}\coloneqq\operatorname{Tr}_{A_{1}}\{\mu^{j}_{B_{1}A_{1}}[\operatorname{\mathds{1}}_{B_{1}}\otimes(\rho^{k}_{A_{1}})^{T}]\}$, and $\tilde{\sigma}^{j,k}_{B_{2}}\coloneqq\operatorname{Tr}_{A_{2}}\{\tilde{\mu}^{j}_{B_{2}A_{2}}[\operatorname{\mathds{1}}_{B_{2}}\otimes(\tilde{\rho}^{k}_{A_{2}})^{T}]\}$. This proves that $d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}$ is a separable state whenever $\mu_{BA^{\prime}}$ and $\rho_{A}$ are separable, therefore condition 2 of Theorem 3 is satisfied. This implies that it is possible to construct a CDRT from the set of separable states which coincides with SEP [37, 38] as noted above. Once again, since the identity matrix is separable, the maximally mixed state is free, as stated in Corollary 4.
§ VIII.  Example: resource theory of NPT entanglement
Looking again at the resource theory of entanglement, one notices that all separable states have positive partial transpose [39, 40]. One can thus consider as free states the bipartite states with positive partial transpose. This set is strictly larger than the set of separable states due to bound entanglement [41], which is of great relevance in quantum information [39, 40, 41, 42, 43, 44, 45]. Since the set of PPT states contains separable states, it automatically satisfies condition 1 of Theorem 3. Let $\mu_{BA^{\prime}}$ and $\rho_{A}$ be free states. We write $(\mu_{B\mathord{:}A^{\prime}}*\rho_{A})^{T_{B_{2}}}$ as
$(\mu_{B\mathord{:}A^{\prime}}*\rho_{A})^{T_{B_{2}}}$ $=\operatorname{Tr}_{A}(\mu_{B\mathord{:}A}^{T_{B_{2}}}(\operatorname{\mathds{1}}_{B}\otimes\rho_{A}^{T}))$ (7)
$=\operatorname{Tr}_{A}[\mu_{BA}^{T_{B_{2}A_{2}}}(\operatorname{\mathds{1}}_{B}\otimes\rho_{A}^{T_{A_{1}}})].$
Since $\rho_{A}$ has positive partial transpose, there exists and eigenbasis {|$\Set{\Ket{\varphi^{j}_{A}}}$ and positive eigenvalues $\lambda_{j}$ such that $\rho_{A}^{T_{A_{1}}}=\sum_{j}\lambda_{j}|\varphi^{j}_{A}\rangle\!\langle\varphi^{j}_{A}|$. This implies that, for all |$\Ket{\psi_{B}}$,
$\Braket{\psi_{B}}{(\mu_{B\mathord{:}A^{\prime}}*\rho_{A})^{T_{B_{2}}}}{\psi_{B}}=$ (8)
$\sum_{j}\lambda_{j}\Braket{\psi_{B}\otimes\varphi_{A}^{j}}{\mu_{BA}^{T_{B_{2}A_{2}}}}{\psi_{B}\otimes\varphi_{A}^{j}}\geq 0,$
where the last inequality follows from the fact that $\mu_{BA}^{T_{B_{2}A_{2}}}\geq 0$, as it is free, and $\lambda_{i}>0$. This is proves that $d_{A}(\mu_{B\mathord{:}A^{\prime}}*\rho_{A})^{T_{B_{2}}}\geq 0$, and that the set of states with positive partial transpose is closed under link product. In this case too, we have shown that the set of free states satisfies the conditions of Theorem 3. The CDRT constructed coincides with the theory of NPT [42, 43, 44, 45].
§ IX.  Conclusions
In this work, we identified a key property of the resource theories of magic, imaginarity, SEP, and NPT, which is that a channel is considered free if and only if its suitably renormalized Choi matrix is a free state. We presented a new way to construct a resource theory from the set of free states based on Choi matrices and we derived which conditions on the free states are necessary and sufficient for constructing a well-defined resource theory. Our interest is driven by the significant advantages that such a property guarantees. These advantages stem from the fact that every question in Choi-defined resource theories can be naturally framed as a question involving free states alone, which often have a well-understood structure. Moreover, we showed that in any CDRT free operations are all CRNG operations. This provides a new construction for CRNG operations when the set of free states satisfy the conditions of Theorem 3. A natural future direction is to explore Choi-defined dynamical resource theories [55, 56, 57], i.e., resource theories for channels, where a superchannel [58, 59] is free if and only its Choi matrix is the Choi matrix of a free channel. It is noteworthy that such property has already been considered in the resource theories for dynamical entanglement [46] and magic [60], where it led to interesting results. The characterization of Choi-defined dynamical resource theories is a particularly challenging problem because the techniques employed in this work cannot be easily translated for superchannels. Indeed, there are some subtleties in the definition of the Choi matrix of a superchannels. For example, in Ref. [58], several different Choi matrices are constructed from the same superchannel, which differ in the order systems are considered. As in the case studied in this Letter, an ordering of systems is mandatory if one wants to construct resource theories from Choi matrices, as inputs and outputs play different roles especially if we want such construction to be generalizable, without ambiguity, to even higher-order maps [61, 62].
Acknowledgements.
The authors are grateful to Matt Wilson for the several in-depth discussions on the subject of this letter and to Gilad Gour for the interesting conversations about this research. E. Z. acknowledges support from the Alberta Graduate Excellence Scholarship (AGES) and from the Eyes High International Doctoral Scholarship. C. M. S. acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) through the Discovery Grant “The power of quantum resources” RGPIN-2022-03025 and the Discovery Launch Supplement DGECR-2022-00119.

### v2

@@
 Abstract
-The resource theories of separable entanglement, non-positive partial transpose entanglement, magic, and imaginarity share an interesting property: An operation is free if and only if its renormalized Choi matrix is a free state.
+Many resource theories share an interesting property: An operation is free if and only if its renormalized Choi matrix is a free state.
 In this letter, we refer to resource theories exhibiting this property as Choi-defined resource theories.
-We demonstrate how and under what conditions one can construct a Choi-defined resource theory, and we prove that when such a construction is possible, the free operations are all and only the completely resource non-generating operations.
+We demonstrate how and under what conditions one can construct a Choi-defined resource theory, and prove that when such a construction is possible, the free operations are all and only the completely resource non-generating operations.
+Moreover, we define resource measures, a complete family of monotones, and conversion distances in such resource theories.
 § I.  Introduction
@@
 These extensions of the resource theory of entanglement are known as the resource theories of separable entanglement (SEP) and non positive partial transpose entanglement (NPT), respectively.
-In other settings, such as in the resource theory of magic [28, 29, 30, 31, 32] and imaginarity [33, 34, 35] the CRNG operations are even more relevant because they coincide with the free operations [31, 32, 35].
-Remarkably, one notices that in each of these instances of resource theories a channel is considered free if and only if its renormalized Choi matrix is a free state [37, 38, 46, 32, 35].
+In other settings, such as in the resource theory of magic [28, 29, 30, 31, 32], imaginarity [33, 34, 35], non-negativity of quantum amplitudes [47], the CRNG operations are even more relevant because they coincide with the free operations.
+Remarkably, one notices that in each of these instances of resource theories a channel is considered free if and only if its renormalized Choi matrix is a free state.
 We call resource theories that exhibit this property Choi-defined resource theories.
-This is a highly valuable property: Every problem associated with free quantum channels can be converted into a problem involving free states, which, in turn, can be conveniently tackled with tools developed in linear algebra, matrix analysis, and convex optimization [37, 38, 32, 35].
+This is a highly valuable property: Every problem associated with free quantum channels can be converted into a problem involving free states, which, in turn, can be conveniently tackled with tools developed in linear algebra, matrix analysis, and convex optimization.
 Given the significance of Choi-defined resource theories and their computational convenience one wonders when it is possible to construct such resource theories from a set of free states.
@@
 Moreover, we show that when such a construction is possible, the resulting set of free operations coincides with the set of CRNG operations associated with those free states.
-These results generalize what has been already observed in the resource theories of magic, imaginarity, NPT, and SEP [37, 38, 46, 31, 32, 35].
+These results provide a constructive definition of CRNG operations and generalize what has been already observed in the resource theories of magic, imaginarity, NPT, SEP, and non-negativity of quantum amplitudes.
+Next, we introduce resource measures, a complete family of monotones, and conversion distances in CDRTs, which can all be computed with conic linear programs (CLP) [48, 49] whenever the set of free states is convex and closed, or even with semidefinite programs (SDP) [50].
+Lastly, we present the resource theory of asymmetry [51], for which the correspondence between the free states and the Choi matrices of the free channels was not already known, and we also characterize which resource theories of asymmetries are CDRTs.
 § II.  Preliminaries
@@
 This may appear to conflict with the condition that the swap channel is a free operation.
-However, in this scenario, the systems in question are bipartite systems [2], i.e., $A=A_{1}A_{2}$ and $B=B_{1}B_{2}$.
-Here the physical separation is between subscript $1$ and $2$.
-The swap channel simply swaps the order of the two bipartite systems, for example, $A_{1}A_{2}B_{1}B_{2}\leftrightarrow B_{1}B_{2}A_{1}A_{2}$.
-If we organize the systems according to the physical separation, the action of the swap channel is $A_{1}B_{1}A_{2}B_{2}\to B_{1}A_{1}B_{2}A_{2}$.
+However, in this scenario, the systems in question are bipartite systems [2], i.e., $S_{1}\coloneqq A_{1}B_{1}$ and $S_{2}\coloneqq A_{2}B_{2}$.
+The physical separation is, as usual, between systems labeled with $A$ and systems labeled with $B$.
+The swap channel simply swaps the order of the two bipartite systems $S_{1}$ and $S_{2}$, i.e., $A_{1}B_{1}A_{2}B_{2}\leftrightarrow A_{2}B_{2}A_{1}B_{1}$.
+If we organize the systems according to the physical separation, the action of the swap channel is $(A_{1}A_{2})(B_{1}B_{2})\to(A_{2}A_{1})(B_{2}B_{1})$.
 Therefore, the two agents are locally exchanging the order of their own quantum systems and there is no exchange of quantum systems between them.
@@
 It is a natural question to characterize when, given a set of free states, the construction of a CDRT is allowed.
-In every well-defined resource theory, the identity is a free operations.
-As a consequence, its renormalized Choi matrix $\frac{1}{d_{A}}\Phi_{AA^{\prime}}$ is free in every CDRT.
-Requesting that resource theories are closed under the sequential composition of free channels (Figure 1), we find a less trivial condition for the set free states.
+In every well-defined resource theory, the identity is a free operation.
+As a consequence, its renormalized Choi state $\frac{1}{d_{A}}\Phi_{AA^{\prime}}$ is free in every CDRT.
+Remark 2.
+The renormalized Choi state is not the maximally entangled state in the resource theories of entanglement, SEP entanglement, or NPT entanglement.
+Indeed, in these resource theories, each system is a pair of pysical systems [2], i.e., $S\coloneqq AB$, and $S^{\prime}\coloneqq A^{\prime}B^{\prime}$ and the spatial separation is between systems $AA^{\prime}$ and $BB^{\prime}$.
+Let {|$\Set{\Ket{a}_{A}}$ and {|$\Set{\Ket{b}_{B}}$ be orthogonal basis for $A$ and $B$, respectively, then
+$\frac{1}{d_{S}}\Phi_{SS^{\prime}}=\frac{1}{d_{A}d_{B}}\sum_{a,b}|ab\rangle\!\langle ab|_{AB}\otimes|ab\rangle\!\langle ab|_{A^{\prime}B^{\prime}}.$ (4)
+If the systems are reorganized according to the pysical separation, the renormalized Choi state becomes
+$\frac{1}{d_{A}d_{B}}\sum_{a}|aa\rangle\!\langle aa|_{AA^{\prime}}\otimes\sum_{b}|bb\rangle\!\langle bb|_{BB^{\prime}}.$ (5)
+This shows that the state $\frac{1}{d_{S}}\Phi_{SS^{\prime}}$ is separable with respect to the spatial separation between $AA^{\prime}$ and $BB^{\prime}$.
+Therefore, it is free in the resource theories mentioned above, so it is not the maximally entangled state on $(AA^{\prime})(BB^{\prime})$.
+Requesting that resource theories be closed under the sequential composition of free channels (Figure 1), we find a less trivial condition for the set free states.
 In Ref.
@@
 Thanks to the link product, we can restate the golden rule of resource theories as a condition on the set of free states.
-That is, the condition that $\mathcal{M}_{A\to B}(\rho_{A})$ be free whenever $\mathcal{M}_{A\to B}$ and $\rho_{A}$ are free becomes $d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}$ is a free state if $\rho_{A}$ and $\mu_{B\mathord{:}A^{\prime}}$ are free states and $\mu_{B\mathord{:}A^{\prime}}$ is the renormalized Choi matrix of a quantum channel.
+That is, the condition that $\mathcal{M}_{A\to B}(\rho_{A})$ be free whenever $\mathcal{M}_{A\to B}$ and $\rho_{A}$ are free becomes that $d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}$ be a free state if $\rho_{A}$ and $\mu_{B\mathord{:}A^{\prime}}$ are free states and $\mu_{B\mathord{:}A^{\prime}}$ is the renormalized Choi matrix of a quantum channel.
 We now state our first main result.
@@
 Notably, this theorem provides an easy construction for CRNG operations when the set of free states satisfies the conditions of Theorem 3.
-For example, the set of separable states satisfies the conditions of Theorem 3, as we show later.
+For example, the set of separable states satisfies the conditions of Theorem 3, as we show in the SM.
 Therefore, Theorems 3 and 5 imply that the CRNG operations are those that have a separable Choi matrix, i.e., the separable operations.
@@
 [56]), but their renormalized Choi matrices are not free states.
-§ V.  Example: resource theory of magic
-Now we give a closer look at the resource theory of magic [28, 29, 30, 31, 32].
-It originates from the observation that it is possible to achieve universal quantum computation by using only Clifford unitaries and some special states called magic states [52].
-Being grounded in quantum computation, systems are usually systems of $n$ qubits [30].
-In this scenario, Clifford unitaries, state preparation and measurements in the computational basis are easy to perform, and therefore free.
-In particular, Clifford unitaries are those transformations that can be implemented with circuits containing only Hadamard, CNOT, and phase-shift gates [53].
-The free operations are called stabilizer operations, and analogously free states are called stabilizer states.
-The resources are magic states and operations.
-We prove now that the set of magic states satisfies the conditions of Theorem 3.
-For the first condition, we denote the number of qubits in $A$ with $n$, and the elements of the computational basis with |$\Ket{x_{1}\dots x_{n}}$, where $x_{j}=0,1$.
-The normalized Choi state |$\Ket{\phi}=\frac{1}{2^{n/2}}\sum_{x_{1}\dots x_{n}}\Ket{x_{1}\dots x_{n}}\Ket{x_{1^{\prime}}\dots x_{n^{\prime}}}$ is obtained by preparing the $n$ entangled states $\frac{1}{\sqrt{2}}\sum_{x_{j}}\Ket{x_{j}x_{j^{\prime}}}$ and by swapping the order of the systems.
-The $j$-th 2-qubit entangled state can be obtained with an Hadamard and a CNOT gate acting on |$\Ket{0_{j}0_{j^{\prime}}}$, and the swap channels can be decomposed into three CNOT gates (see, e.g., Ref.
-[54]).
-All these operations are stabilizer operations, therefore the renormalized Choi state $\frac{1}{d_{A}}\Phi_{AA^{\prime}}$ is a stabilizer state.
-This proves that condition 1 of Theorem 3 is satisfied.
-We now consider the closure under link product of the set of stabilizer states.
-Let $\mu_{BA^{\prime}}$ and $\rho_{A}$ be free states, where $A$ is a $n$-qubit system and $B$ is a $m$-qubit system.
-This implies that there exists a protocol to create $\mu_{BA^{\prime}}\otimes{\rho_{A}}$ for free.
-By Eq.
-(3), the state $d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}$ is obtained by a protocol with postselection:
-  1. 1.
-    Prepare $\mu_{BA^{\prime}}\otimes\rho_{A}$.
-  2. 2.
-    Perform a measurement in the Bell basis for each of the pair of qubits in $A^{\prime}A$
-  3. 3.
-    Postselect on the measurement outcome associated with $\frac{1}{d_{A}}\Phi_{A^{\prime}A}$.
-After the postselection, the system $B$ is in the state $d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}$, as requested.
-Each of the steps of this protocol is a free operation.
-Indeed, the measurement in the Bell basis can be converted into a measurement in the computational basis by adding Hadamard and CNOT gates and these are stabilizer operations.
-Therefore the state $d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}$ is free, as requested by Theorem 3.
-The set of stabilizer states satisfy all the conditions of Theorem 3, and the CDRT constructed form it is the resource theory of magic [32].
-Note that, by Theorem 5, in the resource theory of magic the free operations coincide with the CRNG operations.
-Note that Corollary 4 allows us to predict that the maximally mixed $n$-qubit state $\frac{1}{2^{n}}\operatorname{\mathds{1}}_{2}\otimes\dots\otimes\operatorname{\mathds{1}}_{2}$ is free.
-We can see this directly from the definition of stabilizer states.
-Indeed, each of the qubits is prepared as a probabilistic mixture of elements of the computational basis, and any state preparation in the computational basis is a free operation.
-§ VI.  Example: resource theory of imaginarity
-We now turn our attention to the resource theory of imaginarity [33, 34, 35].
-The aim of this resource theory is to evaluate the role of imaginarity in quantum mechanics.
-A state is free if its density matrix, expressed with respect to a fixed basis, is real.
-States with density matrices containing non-real numbers are resources.
-The Choi state $\frac{1}{d_{A}}\Phi_{AA^{\prime}}$ relative to the fixed basis, has only real coefficients, and therefore is free.
-This proves condition 1 of Theorem 3.
-Condition 2 follows trivially because it can be expressed using only multiplications and traces of real matrices, from which it is possible to obtain only real matrices.
-This proves that it is possible to construct a CDRT from the set of real matrices, and it coincides with the resource theory of imaginarity, where the free operations are all the CRNG operations [35].
-Note that, the maximally mixed state is real too, and therefore free, as predicted by Corollary 4.
-§ VII.  Example: resource theory of separable entanglement
-In the theory of entanglement [5, 6, 7, 8, 9, 10] systems are bipartite systems, and free states are separable states.
-In other words, each of the systems $A$ and $B$ mentioned in Theorem 3 is bipartite, i.e., $A=A_{1}A_{2}$, $B=B_{1}B_{2}$, and the physical separation is between systems with ‘$1$’ as subscript and systems with ‘$2$’ as subscript.
-To prove the first condition of Theorem 3, we fix an ortonormal basis for each system, e.g., {|$\Set{\Ket{a_{1}}}$ is a basis for $A_{1}$, and we write the Choi state as
-$\Phi_{AA^{\prime}}$ $=\sum_{a_{1},a_{2}}|a_{1}a_{2}\rangle\!\langle a^{\prime}_{1}a^{\prime}_{2}|$ (5)
-$=\sum_{a_{1}}|a_{1}\rangle\!\langle a^{\prime}_{1}|\otimes\sum_{a_{2}}|a_{2}\rangle\!\langle a^{\prime}_{2}|.$
-Therefore $\frac{1}{d_{A}}\Phi_{AA^{\prime}}$ is separable, free, and condition 1 of Theorem 3 is satisfied.
-We now prove that the set of separable state is closed under link product.
-Let $\mu_{BA^{\prime}}$ be a separable state, i.e., there exists density matrices $\mu^{j}_{B_{1}A^{\prime}_{1}}$ and $\tilde{\mu}^{j}_{B_{2}A^{\prime}_{2}}$ such that $\mu_{B_{1}A^{\prime}_{1}B_{2}A^{\prime}_{2}}=\sum_{j}p_{j}\mu^{j}_{B_{1}A^{\prime}_{1}}\otimes\tilde{\mu}^{j}_{B_{2}A^{\prime}_{2}}$, where $p_{j}\geq 0$ and $\sum_{j}p_{j}=1$.
-Let $\rho_{A}=\sum_{k}q_{k}\rho^{k}_{A_{1}}\otimes\tilde{\rho}^{k}_{A_{2}}$, where $q_{k}\geq 0$ and $\sum_{k}q_{k}=1$, be another separable state.
-Then
-$\mu_{B\mathord{:}A^{\prime}}*\rho_{A}$ $=\operatorname{Tr}_{A}[\mu_{BA}(\operatorname{\mathds{1}}_{B}\otimes\rho_{A}^{T})]$ (6)
-$=\sum_{j,k}p_{j}q_{k}\sigma^{j,k}_{B_{1}}\otimes\tilde{\sigma}^{j,k}_{B_{2}}.$
-where $\sigma^{j,k}_{B_{1}}\coloneqq\operatorname{Tr}_{A_{1}}\{\mu^{j}_{B_{1}A_{1}}[\operatorname{\mathds{1}}_{B_{1}}\otimes(\rho^{k}_{A_{1}})^{T}]\}$, and $\tilde{\sigma}^{j,k}_{B_{2}}\coloneqq\operatorname{Tr}_{A_{2}}\{\tilde{\mu}^{j}_{B_{2}A_{2}}[\operatorname{\mathds{1}}_{B_{2}}\otimes(\tilde{\rho}^{k}_{A_{2}})^{T}]\}$.
-This proves that $d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}$ is a separable state whenever $\mu_{BA^{\prime}}$ and $\rho_{A}$ are separable, therefore condition 2 of Theorem 3 is satisfied.
-This implies that it is possible to construct a CDRT from the set of separable states which coincides with SEP [37, 38] as noted above.
-Once again, since the identity matrix is separable, the maximally mixed state is free, as stated in Corollary 4.
-§ VIII.  Example: resource theory of NPT entanglement
-Looking again at the resource theory of entanglement, one notices that all separable states have positive partial transpose [39, 40].
-One can thus consider as free states the bipartite states with positive partial transpose.
-This set is strictly larger than the set of separable states due to bound entanglement [41], which is of great relevance in quantum information [39, 40, 41, 42, 43, 44, 45].
-Since the set of PPT states contains separable states, it automatically satisfies condition 1 of Theorem 3.
-Let $\mu_{BA^{\prime}}$ and $\rho_{A}$ be free states.
-We write $(\mu_{B\mathord{:}A^{\prime}}*\rho_{A})^{T_{B_{2}}}$ as
-$(\mu_{B\mathord{:}A^{\prime}}*\rho_{A})^{T_{B_{2}}}$ $=\operatorname{Tr}_{A}(\mu_{B\mathord{:}A}^{T_{B_{2}}}(\operatorname{\mathds{1}}_{B}\otimes\rho_{A}^{T}))$ (7)
-$=\operatorname{Tr}_{A}[\mu_{BA}^{T_{B_{2}A_{2}}}(\operatorname{\mathds{1}}_{B}\otimes\rho_{A}^{T_{A_{1}}})].$
-Since $\rho_{A}$ has positive partial transpose, there exists and eigenbasis {|$\Set{\Ket{\varphi^{j}_{A}}}$ and positive eigenvalues $\lambda_{j}$ such that $\rho_{A}^{T_{A_{1}}}=\sum_{j}\lambda_{j}|\varphi^{j}_{A}\rangle\!\langle\varphi^{j}_{A}|$.
-This implies that, for all |$\Ket{\psi_{B}}$,
-$\Braket{\psi_{B}}{(\mu_{B\mathord{:}A^{\prime}}*\rho_{A})^{T_{B_{2}}}}{\psi_{B}}=$ (8)
-$\sum_{j}\lambda_{j}\Braket{\psi_{B}\otimes\varphi_{A}^{j}}{\mu_{BA}^{T_{B_{2}A_{2}}}}{\psi_{B}\otimes\varphi_{A}^{j}}\geq 0,$
-where the last inequality follows from the fact that $\mu_{BA}^{T_{B_{2}A_{2}}}\geq 0$, as it is free, and $\lambda_{i}>0$.
-This is proves that $d_{A}(\mu_{B\mathord{:}A^{\prime}}*\rho_{A})^{T_{B_{2}}}\geq 0$, and that the set of states with positive partial transpose is closed under link product.
-In this case too, we have shown that the set of free states satisfies the conditions of Theorem 3.
-The CDRT constructed coincides with the theory of NPT [42, 43, 44, 45].
-§ IX.  Conclusions
-In this work, we identified a key property of the resource theories of magic, imaginarity, SEP, and NPT, which is that a channel is considered free if and only if its suitably renormalized Choi matrix is a free state.
-We presented a new way to construct a resource theory from the set of free states based on Choi matrices and we derived which conditions on the free states are necessary and sufficient for constructing a well-defined resource theory.
+§ V.  Quantifying resources in CDRTs
+CDRTs are constructed from a given set of free states that satisfies the conditions of Theorem 3.
+As a consequence, they inherit all the resource measures that depend only on the set of free states.
+Examples of such resource measures are the robustness, the generalized robustness, and all the ones based on quantum divergences (see Ref.
+[1] for a review).
+Among these, the max-relative entropy [57] is of particular interest in the context of CDRTs.
+It is defined as
+$D_{\max}(\rho_{A}\|\sigma_{A})=\log\min\Set{\lambda\mid\rho_{A}\leq\lambda\sigma_{A}},$ (7)
+when $\operatorname{supp}\rho_{A}\subseteq\operatorname{supp}\sigma_{A}$ and +$\infty$+ ∞ otherwise. The extension of the max-relative entropy to quantum channels [58, 59] satisfies [60]
+$D_{\max}(\mathcal{M}_{A\to B}\|\mathcal{N}_{A\to B})=D_{\max}(\mu_{BA}\|\nu_{BA}),$ (8)
+where, $\mu_{BA}$ and $\nu_{BA}$ are the renormalized Choi matrices of $\mathcal{M}_{A\to B}$ and $\mathcal{N}_{A\to B}$, respectively.
+In a CDRT, this property allows us to define resource measures for states [1] and channels [61] that depend uniquely on the set of free states:
+$D_{\max}^{\mathfrak{F}}(\rho_{A})=\inf_{\omega_{A}\in\mathfrak{F}(A)}D_{\max}(\rho_{A}\|\omega_{A}),$ (9)
+$D_{\max}^{\mathfrak{F}}(\mathcal{M}_{A\to B})=\inf_{\begin{subarray}{c}\omega_{BA}\in\mathfrak{F}(BA),\\ d_{A}\operatorname{Tr}_{B}\omega_{BA}=\operatorname{\mathds{1}}_{A}\end{subarray}}D_{\max}(\mu_{BA}\|\omega_{BA}),$
+where $\mathfrak{F}(A)$ denotes the set of free states of system $A$.
+Both these resource measures are finite in every CDRTs because for every $\rho_{A}$, $\operatorname{supp}\rho_{A}\subseteq\operatorname{supp}\operatorname{\mathds{1}}_{A}$ and $\frac{1}{d_{A}}\operatorname{\mathds{1}}_{A}$ is free.
+The same is not true for all resource theories.
+Moreover, these resource measures can be computed with a CLP whenever the set of free states is convex and closed, and with a SDP whenever the membership problem for the set of free states is an SDP (see the SM).
+As a direct consequence of the results in Ref. [60], the resource measures in Eq. (9) satisfy
+$D_{\max}^{\mathfrak{F}}(\mathcal{M}_{A\to B}(\rho_{A}))\leq D_{\max}^{\mathfrak{F}}(\rho_{A})+D_{\max}^{\mathfrak{F}}(\mathcal{M}_{A\to B})$ (10)
+for all states $\rho_{A}$ and channels $\mathcal{M}_{A\to B}$.
+For the reasons mentioned above, this inequality, which is valid for all resource theories, is particularly meaningful in the context of CDRTs.
+When we interpret $D^{\mathfrak{F}}$ as the value of a resource, it implies that value of the output of a quantum operation applied to a quantum state is always less than the sum of the values of the operation and the input state.
+Equivalently, for every $\rho_{A}$, $\sigma_{B}$, and $\mathcal{M}_{A\to B}$ such that $\sigma_{B}=\mathcal{M}_{A\to B}(\rho)$, the cost of preparing $\sigma_{B}$ is always less or equal than the cost of preparing $\rho_{A}$ and then converting it to $\sigma_{B}$ with the channel $\mathcal{M}_{A\to B}$.
+An inequality similar to the one in Eq.
+(10) was derived in Ref.
+[62].
+The divergences that appear in that inequality are the relative entropy and the regularized channel relative entropy.
+Instead, in our case, only the max-relative entropy is used, and no regularization is needed.
+§ VI.  Resource conversion in CDRTs
+Where the formalism of CDRTs provides the greatest insight is in all quantities that are connected with resource conversion or manipulation via free operations.
+Indeed at the core of every resource theory there is a preorder of resources defined as $\rho_{A}\succ\sigma_{B}$ if there exists a free channel $\mathcal{M}_{A\to B}$ such that $\sigma_{B}=\mathcal{M}_{A\to B}(\rho_{A})$.
+In a CDRT, this condition is equivalent to requiring the existence of a free states $\mu_{BA}$, such that $\operatorname{Tr}_{B}\mu_{BA}=\frac{1}{d_{A}}\operatorname{\mathds{1}}_{A}$ and $\sigma_{B}=d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}$.
+As a consequence, many quantities or results that involve free channels can be redefined in terms of the link product of free states.
+A first example is the complete family of monotones defined in Ref.
+[61] for convex and closed resource theories, which in the case of a convex and closed CDRT can be expressed as
+$f_{\tau_{B}}(\rho_{A})=\max_{\begin{subarray}{c}\omega_{BA}\in\mathfrak{F}(BA),\\ d_{A}\operatorname{Tr}_{B}\omega_{BA}=\operatorname{\mathds{1}}_{A}\end{subarray}}\operatorname{Tr}_{BA}[\omega_{BA}(\tau_{B}\otimes\rho_{A}^{T})].$ (11)
+Being a complete family of monotones, it satisfies $f_{\tau_{B}}(\rho_{A})\geq f_{\tau_{B}}(\sigma_{B})$ for all density matrices $\tau_{B}$ if and only if $\rho_{A}$ can be converted to $\sigma_{B}$ with a free operation.
+The advantage of this approach is that the optimization is over states and not channels.
+In particular, the quantity above can be computed with a CLP (see the SM).
+As a second example, we present conversion distances in CDRTs.
+These distances estimate how close to a target state an initial state can be converted by using only free operations.
+In a CDRT, they satisfy
+$D(\rho_{A}\to\sigma_{B})$ $=\inf_{\mathcal{M}_{A\to B}\in\mathfrak{F}(A\to B)}D(\sigma_{B},\mathcal{M}_{A\to B}(\rho_{A}))$ (12)
+$=\inf_{\begin{subarray}{c}\mu_{BA}\in\mathfrak{F}(BA),\\ d_{A}\operatorname{Tr}_{B}\mu_{BA}=\operatorname{\mathds{1}}_{A}\end{subarray}}D(\sigma_{B},d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}),$
+where $D(\cdot,\cdot)$ is a function on quantum states that satisfies the data processing inequality [1].
+Once again, the minimization is only over states and this can simplify the computation of the conversion distances.
+In particular, many metrics, such as the trace distance or the purified distance, can be computed with SDPs [50].
+Consequently, the conversion distances based on such metrics are SDPs too whenever the membership problem for the set of free states can be expressed as an SDP, e.g., in the case of the resource theories of imaginarity and PPT.
+§ VII.  Examples of CDRTs
+As mentioned in the introduction, known results about the resource theories of SEP, PPT, magic, imaginarity, and non-negative quantum amplitudes imply that these resource theories are CDRTs and we show in the SM that their sets of free states satisfy the conditions of Theorem 3.
+Of more interest are those resource theories where the identification of free operations with CD operation has not been already established.
+For example, to the best of our knowledge, there is no result yet in this direction for the resource theory of asymmetry.
+In this section, we derive under what conditions resource theories of asymmetry [51, 14] are CDRTs.
+In the resource theory of asymmetry, two parties, Alice and Adam have access to the same quantum systems but they do not share a reference frame.
+The relation between reference frames is described by an element $g$ of a compact group $G$ identified by the symmetries of the problem.
+If Alice and Adam do not know how their reference frames are related, then Alice can prepare only states for Adam that are $G$-invariant, that is:
+$\rho_{A}\,\text{is free}\,\Leftrightarrow\mathcal{U}_{g}^{A}(\rho_{A})=\rho_{A},\,\forall g\in G,$ (13)
+where $\mathcal{U}_{g}^{A}(\rho_{A})\coloneq U^{A}_{g}\rho_{A}(U^{A}_{g})^{\dagger}$, and $U_{g}$ is the unitary representation of $g$ on $A$.
+Similarly, if Adam asks Alice to perform a deterministic quantum operation, then she can only perform $G$-covariant channels:
+$\mathcal{M}_{A\to B}\,\text{is free}\,\Leftrightarrow\mathcal{U}^{B}_{g}\circ\mathcal{M}_{A\to B}\circ(\mathcal{U}_{g}^{A})^{\dagger}=\mathcal{M}_{A\to B}$ (14)
+for all $g\in G$.
+Every different finite or compact Lie group $G$ generates a different resource theory, and the following proposition characterizes the ones that are CDRTs.
+Proposition 6.
+A resource theory of asymmetry is a CDRT if and only if the renormalized Choi state is free.
+We point out that this is equivalent to asking that all the unitary representations are real (see the SM for details).
+An example of such resource theory is the parity resource theory generated by the group $\mathbb{Z}_{2}$ [51].
+§ VIII.  Conclusions
+In this work, we identified a key property of the resource theories of magic, imaginarity, SEP, NPT, and non-negativity of quantum amplitudes, which is that a channel is considered free if and only if its suitably renormalized Choi matrix is a free state.
+We presented a new way to construct a resource theory from the set of free states based on the Choi isomorphism and we derived which conditions on the free states are necessary and sufficient for constructing a mathematically consistent Choi-defined resource theory.
 Our interest is driven by the significant advantages that such a property guarantees.
 These advantages stem from the fact that every question in Choi-defined resource theories can be naturally framed as a question involving free states alone, which often have a well-understood structure.
+Indeed, we introduce several resource quantifiers which can all be computed with CLPs if the set of free states is convex and closed.
 Moreover, we showed that in any CDRT free operations are all CRNG operations.
-This provides a new construction for CRNG operations when the set of free states satisfy the conditions of Theorem 3.
+This provides a new constructive way to define CRNG operations when the set of free states satisfy the conditions of Theorem 3.
 A natural future direction is to explore Choi-defined dynamical resource theories [61, 63, 64], i.e., resource theories for channels, where a superchannel [65, 66] is free if and only its Choi matrix is the Choi matrix of a free channel.
@@
 Acknowledgements.
-The authors are grateful to Matt Wilson for the several in-depth discussions on the subject of this letter and to Gilad Gour for the interesting conversations about this research.
+The authors are grateful to Matt Wilson for the several in-depth discussions on the subject of this letter and to Gilad Gour and Nathaniel Johnston for the interesting conversations about this research.
 E.

### v3

@@
 In this letter, we refer to resource theories exhibiting this property as Choi-defined resource theories.
-We demonstrate how and under what conditions one can construct a Choi-defined resource theory, and prove that when such a construction is possible, the free operations are all and only the completely resource non-generating operations.
-Moreover, we define resource measures, a complete family of monotones, and conversion distances in such resource theories.
+We demonstrate how and under what conditions one can construct a Choi-defined resource theory, and we prove that when such a construction is possible, the free operations are all and only the completely resource non-generating operations.
+Moreover, we examine resource measures, a complete family of monotones, and conversion distances in such resource theories.
 § I.  Introduction
@@
 The new technological developments, which are now becoming commercially available, demand a rigorous and thorough understanding of quantum resources, in order to quantify them precisely, and optimize their use.
-This is accomplished with the framework of resource theories [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35].
+This is accomplished with the framework of resource theories [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36].
 Resource theories are based on the simple observation that agents in a laboratory can only perform a subset of all possible deterministic quantum operations due to the various constraints on the agents and the systems involved.
@@
 We call resource theories that exhibit this property Choi-defined resource theories.
-This is a highly valuable property: Every problem associated with free quantum channels can be converted into a problem involving free states, which, in turn, can be conveniently tackled with tools developed in linear algebra, matrix analysis, and convex optimization.
+This is a highly valuable property: Every problem associated with free quantum channels can be converted into a problem involving free states.
 Given the significance of Choi-defined resource theories and their computational convenience one wonders when it is possible to construct such resource theories from a set of free states.
@@
 Next, we introduce resource measures, a complete family of monotones, and conversion distances in CDRTs, which can all be computed with conic linear programs (CLP) [36, 49] whenever the set of free states is convex and closed, or even with semidefinite programs (SDP) [50].
-Lastly, we present the resource theory of asymmetry [51], for which the correspondence between the free states and the Choi matrices of the free channels was not already known, and we also characterize which resource theories of asymmetries are CDRTs.
 § II.  Preliminaries
@@
   5. 5.
-    parallel composition of free channels is free
+    parallel composition of free channels is free.
 In other words, an agent is allowed to do nothing on a system, to change the order in which they describe systems, and to discard a system.
@@
 The renormalized Choi state is not the maximally entangled state in the resource theories of entanglement, SEP entanglement, or NPT entanglement.
-Indeed, in these resource theories, each system is a pair of pysical systems [2], i.e., $S\coloneqq AB$, and $S^{\prime}\coloneqq A^{\prime}B^{\prime}$ and the spatial separation is between systems $AA^{\prime}$ and $BB^{\prime}$.
+Indeed, in these resource theories, each system is a pair of physical systems [2], i.e., $S\coloneqq AB$, and $S^{\prime}\coloneqq A^{\prime}B^{\prime}$ and the spatial separation is between systems $AA^{\prime}$ and $BB^{\prime}$.
 Let {|$\Set{\Ket{a}_{A}}$ and {|$\Set{\Ket{b}_{B}}$ be orthogonal basis for $A$ and $B$, respectively, then
-$\frac{1}{d_{S}}\Phi_{SS^{\prime}}=\frac{1}{d_{A}d_{B}}\sum_{a,b}|ab\rangle\!\langle ab|_{AB}\otimes|ab\rangle\!\langle ab|_{A^{\prime}B^{\prime}}.$ (4)
+$\frac{1}{d_{S}}\Phi_{SS^{\prime}}=\frac{1}{d_{A}d_{B}}\sum_{a,b,\tilde{a},\tilde{b}}|ab\rangle\!\langle\tilde{a}\tilde{b}|_{AB}\otimes|ab\rangle\!\langle\tilde{a}\tilde{b}|_{A^{\prime}B^{\prime}}.$ (4)
 If the systems are reorganized according to the pysical separation, the renormalized Choi state becomes
-$\frac{1}{d_{A}d_{B}}\sum_{a}|aa\rangle\!\langle aa|_{AA^{\prime}}\otimes\sum_{b}|bb\rangle\!\langle bb|_{BB^{\prime}}.$ (5)
+$\frac{1}{d_{A}d_{B}}\sum_{a,\tilde{a}}|aa\rangle\!\langle\tilde{a}\tilde{a}|_{AA^{\prime}}\otimes\sum_{b,\tilde{b}}|bb\rangle\!\langle\tilde{b}\tilde{b}|_{BB^{\prime}}$ (5)
+$=\frac{1}{d_{A}}\Phi_{AA^{\prime}}\otimes\frac{1}{d_{B}}\Phi_{BB^{\prime}}.$
 This shows that the state $\frac{1}{d_{S}}\Phi_{SS^{\prime}}$ is separable with respect to the spatial separation between $AA^{\prime}$ and $BB^{\prime}$.
@@
 [55]), but their renormalized Choi matrices are not free states.
-§ V.  Quantifying resources in CDRTs
+§ V.  Examples of CDRTs
+As mentioned in the introduction, known results about the resource theories of SEP, PPT, magic, imaginarity, and non-negative quantum amplitudes imply that these resource theories are CDRTs and we show in the SM that their sets of free states satisfy the conditions of Theorem 3.
+The attempt to construct a CDRT from $k$-extendible states made in Ref.
+[56] did not result in a proper resource theory because $k$-extendible states do not satisfy condition 2 of Theorem 3.
+The resource theory of $k$-unextendibility was later defined using different free operations [57].
+Of more interest are those resource theories where the identification of free operations with CD operation has not been already established.
+For example, to the best of our knowledge, there is no explicit result yet in this direction for the resource theory of asymmetry [58, 14].
+In the SM, we show that a resource theory of asymmetry is a CDRT if and only if the renormalized Choi state is free, which is equivalent to asking that all the unitary representations are real (cf.
+Supplementary Information of Ref.
+[59]).
+§ VI.  Optimization problems in CDRTs
+In CDRTs, every optimization problem over the set of free channels can be converted into an optimization problem over the set of free states. That is,
+$\max_{\mathcal{M}_{A\to B}\in\mathfrak{F}(A\to B)}f(\mathcal{M}_{A\to B})$ (7)
+becomes
+$\max_{\begin{subarray}{c}\mu_{BA^{\prime}}\in\mathfrak{F}(BA^{\prime}),\\ d_{A^{\prime}}\operatorname{Tr}_{B}\mu_{BA^{\prime}}=\operatorname{\mathds{1}}_{A^{\prime}}\end{subarray}}\tilde{f}(\mu_{BA^{\prime}}).$ (8)
+Here, $f$ is a function from the set of quantum channels to the $\mathbb{R}$, $\mathfrak{F}(A\to B)$ and $\mathfrak{F}(BA^{\prime})$ are the sets of free channels from $A$ to $B$ and free states on $BA^{\prime}$, and $\tilde{f}$ is the composition of $f$ with the inverse Choi isomorphism defined in Eq.
+(2), and it satisfies $\tilde{f}(\mu_{BA^{\prime}})=f(\mathcal{M}_{A\to B})$ whenever $\mu_{BA^{\prime}}$ is the renormalized Choi matrix of $\mathcal{M}_{A\to B}$.
+Moreover, if the set of free states is convex and $\tilde{f}$ is linear, the optimization problem can be expressed as a CLP (see SM for details) or even as an SDP.
+If one has a set of free states that does not satisfy the conditions of Theorem 3, one can still consider the CRNG operations.
+However, in this case, there is no universal way of expressing the constraints in optimization problems as a single constraint in terms of free states.
+§ VII.  Quantifying resources in CDRTs
 CDRTs are constructed from a given set of free states that satisfies the conditions of Theorem 3.
@@
 $D_{\max}(\rho_{A}\|\sigma_{A})=\log\min\Set{\lambda\mid\rho_{A}\leq\lambda\sigma_{A}},$ (9)
-when $\operatorname{supp}\rho_{A}\subseteq\operatorname{supp}\sigma_{A}$ and +$\infty$+ ∞ otherwise. The extension of the max-relative entropy to quantum channels [58, 59] satisfies [60]
-$D_{\max}(\mathcal{M}_{A\to B}\|\mathcal{N}_{A\to B})=D_{\max}(\mu_{BA}\|\nu_{BA}),$ (8)
-where, $\mu_{BA}$ and $\nu_{BA}$ are the renormalized Choi matrices of $\mathcal{M}_{A\to B}$ and $\mathcal{N}_{A\to B}$, respectively.
-In a CDRT, this property allows us to define resource measures for states [1] and channels [61] that depend uniquely on the set of free states:
-$D_{\max}^{\mathfrak{F}}(\rho_{A})=\inf_{\omega_{A}\in\mathfrak{F}(A)}D_{\max}(\rho_{A}\|\omega_{A}),$ (9)
-$D_{\max}^{\mathfrak{F}}(\mathcal{M}_{A\to B})=\inf_{\begin{subarray}{c}\omega_{BA}\in\mathfrak{F}(BA),\\ d_{A}\operatorname{Tr}_{B}\omega_{BA}=\operatorname{\mathds{1}}_{A}\end{subarray}}D_{\max}(\mu_{BA}\|\omega_{BA}),$
-where $\mathfrak{F}(A)$ denotes the set of free states of system $A$.
-Both these resource measures are finite in every CDRTs because for every $\rho_{A}$, $\operatorname{supp}\rho_{A}\subseteq\operatorname{supp}\operatorname{\mathds{1}}_{A}$ and $\frac{1}{d_{A}}\operatorname{\mathds{1}}_{A}$ is free.
+when $\operatorname{supp}\rho_{A}\subseteq\operatorname{supp}\sigma_{A}$ and +$\infty$+ ∞ otherwise. The resource measure based on the max-relative entropy is
+$D_{\max}^{\mathfrak{F}}(\rho_{A})=\inf_{\omega_{A}\in\mathfrak{F}(A)}D_{\max}(\rho_{A}\|\omega_{A}).$ (10)
+The extension of these quantities to quantum channels are [61, 62]
+$D_{\max}(\mathcal{M}_{A\to B}\|\mathcal{N}_{A\to B})=$ (11)
+$\quad\sup_{\rho_{RA}}D_{\max}(\mathcal{I}_{R}\otimes\mathcal{M}_{A\to B}(\rho_{RA})\|\mathcal{I}_{R}\otimes\mathcal{N}_{A\to B}(\rho_{RA})),$
+$D_{\max}^{\mathfrak{F}}(\mathcal{M}_{A\to B})=\inf_{\mathcal{N}_{A\to B}\in\mathfrak{F}(A\to B)}D_{\max}(\mathcal{M}_{A\to B}\|\mathcal{N}_{A\to B}).$
+Notably, the max-relative entropy for quantum channels satisfies [63]
+$D_{\max}(\mathcal{M}_{A\to B}\|\mathcal{N}_{A\to B})=D_{\max}(\mu_{BA^{\prime}}\|\nu_{BA^{\prime}}),$ (12)
+where $\mu_{BA^{\prime}}$ and $\nu_{BA^{\prime}}$ are the renormalized Choi matrices of $\mathcal{M}_{A\to B}$ and $\mathcal{N}_{A\to B}$, respectively.
+In a CDRT, this property is particularly relevant because it allows us to write the optimization problem for $D_{\max}^{\mathfrak{F}}(\mathcal{M}_{A\to B})$ as
+$D_{\max}^{\mathfrak{F}}(\mathcal{M}_{A\to B})=\inf_{\begin{subarray}{c}\omega_{BA^{\prime}}\in\mathfrak{F}(BA^{\prime}),\\ d_{A^{\prime}}\operatorname{Tr}_{B}\omega_{BA^{\prime}}=\operatorname{\mathds{1}}_{A^{\prime}}\end{subarray}}D_{\max}(\mu_{BA^{\prime}}\|\omega_{BA^{\prime}}).$ (13)
+Notice that, in every CDRT, $D_{\max}^{\mathfrak{F}}$ is finite for both states and channels because the set of free states always contain a full rank state, see Corollary 4.
 The same is not true for all resource theories.
-Moreover, these resource measures can be computed with a CLP whenever the set of free states is convex and closed, and with a SDP whenever the membership problem for the set of free states is an SDP (see the SM).
-As a direct consequence of the results in Ref. [60], the resource measures in Eq. (9) satisfy
+As a direct consequence of the results in Ref. [63], the resource measures based on the max-relative entropy satisfy
 $D_{\max}^{\mathfrak{F}}(\mathcal{M}_{A\to B}(\rho_{A}))\leq D_{\max}^{\mathfrak{F}}(\rho_{A})+D_{\max}^{\mathfrak{F}}(\mathcal{M}_{A\to B})$ (14)
@@
 An inequality similar to the one in Eq.
-(10) was derived in Ref.
+(14) based on the quantum relative entropy and its regularization was derived in Ref.
 [64].
-The divergences that appear in that inequality are the relative entropy and the regularized channel relative entropy.
-Instead, in our case, only the max-relative entropy is used, and no regularization is needed.
-§ VI.  Resource conversion in CDRTs
+§ VIII.  Resource conversion in CDRTs
 Where the formalism of CDRTs provides the greatest insight is in all quantities that are connected with resource conversion or manipulation via free operations.
@@
 Being a complete family of monotones, it satisfies $f_{\tau_{B}}(\rho_{A})\geq f_{\tau_{B}}(\sigma_{B})$ for all density matrices $\tau_{B}$ if and only if $\rho_{A}$ can be converted to $\sigma_{B}$ with a free operation.
-The advantage of this approach is that the optimization is over states and not channels.
-In particular, the quantity above can be computed with a CLP (see the SM).
 As a second example, we present conversion distances in CDRTs.
@@
 $D(\rho_{A}\to\sigma_{B})$ $=\inf_{\mathcal{M}_{A\to B}\in\mathfrak{F}(A\to B)}D(\sigma_{B},\mathcal{M}_{A\to B}(\rho_{A}))$ (16)
-$=\inf_{\begin{subarray}{c}\mu_{BA}\in\mathfrak{F}(BA),\\ d_{A}\operatorname{Tr}_{B}\mu_{BA}=\operatorname{\mathds{1}}_{A}\end{subarray}}D(\sigma_{B},d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}),$
+$=\inf_{\begin{subarray}{c}\mu_{BA^{\prime}}\in\mathfrak{F}(BA^{\prime}),\\ d_{A^{\prime}}\operatorname{Tr}_{B}\mu_{BA^{\prime}}=\operatorname{\mathds{1}}_{A^{\prime}}\end{subarray}}D(\sigma_{B},d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}),$
 where $D(\cdot,\cdot)$ is a function on quantum states that satisfies the data processing inequality [1].
-Once again, the minimization is only over states and this can simplify the computation of the conversion distances.
-In particular, many metrics, such as the trace distance or the purified distance, can be computed with SDPs [50].
-Consequently, the conversion distances based on such metrics are SDPs too whenever the membership problem for the set of free states can be expressed as an SDP, e.g., in the case of the resource theories of imaginarity and PPT.
-§ VII.  Examples of CDRTs
-As mentioned in the introduction, known results about the resource theories of SEP, PPT, magic, imaginarity, and non-negative quantum amplitudes imply that these resource theories are CDRTs and we show in the SM that their sets of free states satisfy the conditions of Theorem 3.
-Of more interest are those resource theories where the identification of free operations with CD operation has not been already established.
-For example, to the best of our knowledge, there is no result yet in this direction for the resource theory of asymmetry.
-In this section, we derive under what conditions resource theories of asymmetry [51, 14] are CDRTs.
-In the resource theory of asymmetry, two parties, Alice and Adam have access to the same quantum systems but they do not share a reference frame.
-The relation between reference frames is described by an element $g$ of a compact group $G$ identified by the symmetries of the problem.
-If Alice and Adam do not know how their reference frames are related, then Alice can prepare only states for Adam that are $G$-invariant, that is:
-$\rho_{A}\,\text{is free}\,\Leftrightarrow\mathcal{U}_{g}^{A}(\rho_{A})=\rho_{A},\,\forall g\in G,$ (13)
-where $\mathcal{U}_{g}^{A}(\rho_{A})\coloneq U^{A}_{g}\rho_{A}(U^{A}_{g})^{\dagger}$, and $U_{g}$ is the unitary representation of $g$ on $A$.
-Similarly, if Adam asks Alice to perform a deterministic quantum operation, then she can only perform $G$-covariant channels:
-$\mathcal{M}_{A\to B}\,\text{is free}\,\Leftrightarrow\mathcal{U}^{B}_{g}\circ\mathcal{M}_{A\to B}\circ(\mathcal{U}_{g}^{A})^{\dagger}=\mathcal{M}_{A\to B}$ (14)
-for all $g\in G$.
-Every different finite or compact Lie group $G$ generates a different resource theory, and the following proposition characterizes the ones that are CDRTs.
-Proposition 6.
-A resource theory of asymmetry is a CDRT if and only if the renormalized Choi state is free.
-We point out that this is equivalent to asking that all the unitary representations are real (see the SM for details).
-An example of such resource theory is the parity resource theory generated by the group $\mathbb{Z}_{2}$ [51].
-§ VIII.  Conclusions
+§ IX.  Conclusions
 In this work, we identified a key property of the resource theories of magic, imaginarity, SEP, NPT, and non-negativity of quantum amplitudes, which is that a channel is considered free if and only if its suitably renormalized Choi matrix is a free state.
@@
 Moreover, we showed that in any CDRT free operations are all CRNG operations.
-This provides a new constructive way to define CRNG operations when the set of free states satisfy the conditions of Theorem 3.
+This provides a new constructive way to define CRNG operations when the set of free states satisfies the conditions of Theorem 3.
 A natural future direction is to explore Choi-defined dynamical resource theories [65, 66, 67], i.e., resource theories for channels, where a superchannel [68, 69] is free if and only its Choi matrix is the Choi matrix of a free channel.
@@
 Z.
-acknowledges support from the Alberta Graduate Excellence Scholarship (AGES) and from the Eyes High International Doctoral Scholarship.
+acknowledges support from the Eric Milner’s Graduate Scholarship, the Alberta Graduate Excellence Scholarship (AGES) and from the Eyes High International Doctoral Scholarship.
 C.